Use FOIL to find the product (2b+9)(5b+7)
Answer by Theo(13342)

FOIL is First Outer Inner Last
This means:
Multiply the first terms in each factor first.
Multiply the outer terms in each factor next.
Multiply the inner terms in each factor next.
Multiply the last terms in each factor next.
Then you add all the results of each multiplication together and you combine like terms and you have your answer.
Your problem is:
(2b+9)*(5b+7)
Multiply the First terms in each factor first.
That would be 2b * 5b which equals 10b^2
Multiply the Outer terms in each factor next.
That would be 2b * 7 = 14b
Multiply the Inner terms in each factor next.
That would be 9 * 5b = 45b
Multiply the Outer terms in each factor next.
That would be 9 * 7 = 63
Add all results of each multiplication together and you get:
10b^2 + 14b + 45b + 63
Combine like terms and you get:
10b^2 + 59b + 63
That's your answer.